Determine the density of oxygen at 150,000 Pa pressure and Temperature = 27°C. State your answer to nearest 0.001 kg/m3, follow the standard rounding rules in stating answer, and do not attempt to write the units. Take the molar mass of molecular oxygen to be 32 g/mol, R=8.31 JK-1mol-1 and take the absolute zero to be -273°C.
1
Expert's answer
2019-09-24T09:07:41-0400

P

"PV=vRT ;\rho=m/V;v=m/M\implies"

"MP=\rho RT;\rho =\frac{MP}{RT}=\frac{32·10^{-3}·150000}{8.31·(273+27)}=" 1.925

Answer 1.925